8-pack control system working group Meeting Notes

30 September 2003


In attendance: R. Chestnut, K. Kim, J. Nelson, J. Rock, M. Ross, S. Smith, S. Tantawi

Note taker/typer: J. Nelson



Status Items:


      MacNair Digitizer Support

            Kukhee is almost done.  Maybe next week we can start testing.


      Trickle Down of Sled and Phase Numbers

            Coming soon.  Kukhee will start working on this after the MacNair digitizer support is finished.  He hopes this will be finished by the time we start testing.



            Judy reported that database work is done.  There’s still about half a dozen panels left to do with a few hundred PVs.  MCC Op Matt Boyes is helping out with this.


      AB Purple Status

            The EPICS timeout was changed a week or so ago.  Judy will check cmlog to see if the frequency of the purple status has changed.


      SLED/LD Databases/Panel Mods (incl plots)

            Judy repoted that all displays are ready.  Janice will scope them out to make sure the specifications were sufficient, then we’re ready to test in October.

            Steve said the LLRF enable buttons can be removed from the epics displays.


      Hardware Deadman

            We believe that the software support for this is in place.  If the logic is not: deadman DIO bit is toggled at the processing rate (~60Hz), Ron with email Steve and Joe.

            Joe and Steve will work on getting the hardware aspect working.


      IQAs and PHITs

            Steve reported we have 4 IQAs with another coming soon.  We need 6 for this phase of the project (loads on the roof)

            When RF folks decide to instrument the load reflected coupler ports (not just the forward as is the case now), Steve will let SWE know which ADC/SIS/IQA channels will be used for them.

            PHITs are ready.


      Peak Power Meters

            Folks couldn’t remember how many heads we have out there, were some sent for repair?  Janice will ask NLCTA OPs to see if they remember.

            There are 4 PPM channels on the dance floor now.

            LLRF folks will bring a PPM during the testing


      Knobs for Controlling SLED Movers

            Janice will work with Sami to make a sum and a difference knob for these.  (Janice also promised to show him how to get to the 8pack processing loop panel.)



Things to Test (tentative start date is 10/13)

      SLED/Phase trickle down

      Processing loop with real values

      ADC readback & pwr displays

      SIS readback (I&Q, A&P, calibrated waveforms, etc)

      (longer term) Calibration (LLRF) (matlab routines??)

      RE/vac/proc loop/etc trips, also diode SW reset

      Check and test that all interesting channels are being archived appropriately.